OpenStudy (anonymous):

One number is 4 more than another, and their sum is 60. What is the smaller number? 26 28 32

OpenStudy (anonymous):

28?

OpenStudy (solomonzelman):

4a=b a+b=60

OpenStudy (solomonzelman):

You need to sovle for \(a\).

OpenStudy (anonymous):

More than makes me think of addition, not multiplication /:

OpenStudy (anonymous):

9 is 4 more than 5. See how that sounds right?

OpenStudy (solomonzelman):

oh, yeah ... also based on the options at least this is correct.

OpenStudy (anonymous):

wait so its 26?

OpenStudy (anonymous):

\[\large \sf x=y+4\]\[\large \sf x+y=60\]\[\large \sf (y+4)+y=60\]

OpenStudy (anonymous):

You following so far?

OpenStudy (anonymous):

(It's not 26)